how is using one ,two or three variables to equations different from each other and how are they the same?
Answer by Nate(3500)

Equations using one variable only need one equation to solve.
x + 1 = 2
x = 1
Equations using two variables generally needs two equations to solve.
x + y = 2
x - y = 0
x = 1
y = 1
Equations using three variables generally needs three equations to solve.
x + z + 2y = 4
x + z - y = 1
2x + 2z - 4y = 0
x = 1
z = 1
y = 1
You can obviously find many reasons to described how they are the same.